# Classical Benchmarks for Variational Quantum Eigensolver Simulations of the Hubbard Model

### v1

@@ II.2.3 Unitary coupled cluster (UCC) ansatz
III Results
@@
We now systematically study the impact of varying the number of variational parameters and the ansatz, the strength of the electronic interactions, and the strength of disorder and off-site Coulomb interactions, on the quality of the VQE solutions of the Hubbard model with respect to DMRG. Moreover, we benchmark the performance of using an overlap-based optimization, compared to the more conventional energy-based optimization. The optimized VQE energies for all cases presented below are given in the Supplemetary Material.
III.1 Impact of the ansatz and the number of variational parameters
@@
We start our analysis by describing the VQE result for the Hubbard model of different system sizes in $d=1$ and $d=2$ square lattice geometries. We employ open boundary conditions throughout. We will first explore how the NP ansatz, which was designed specifically for solving the Hubbard model on quantum hardware, performs as a function of lattice size and number of ansatz parameters. Our benchmarks make use of PQCs of up to thirteen layers, with each layer contributing $10N_{x}N_{y}-4N_{x}-4N_{y}$ parameters.
@@
In Fig. 1a we plot for several one-dimensional lattices the difference between our VQE energy and the DMRG benchmark per lattice site, for $U/t=2$. This interaction strength lies in the weakly- to moderately-correlated regime [55, 32]. As our ground state energy from VQE we plot our best estimate, obtained as the lowest value from ten independent optimizations with random starting points. The shaded regions in Fig. 1a indicate the full range of energies obtained over all optimizations. While it is clear that the VQE optimization converges to the ground state energy even for the larger lattices as more parameters are added to the PQC, the convergence becomes increasingly slow for larger lattice sizes, and even plateaus to errors of order $10^{-2}$ in those cases. For the $1\times 12$ lattice with $13$ layers of the NP ansatz we have also performed an extended optimization, where we performed a hundred independent optimizations (rather than ten) starting from different random parameters. Nevertheless, even when we explore a larger part of phase space, we do not find any appreciable reduction in the ground state energy produced by VQE. Fig. 1b visualizes the performance of VQE on a logarithmic scale for two-dimensional lattices at $U/t=2$. We again see that the convergence of VQE using the NP ansatz slows down with increasing system size, consistent with our benchmarks of the one-dimensional case.
@@
To further understand the convergence behavior of the VQE energy towards the DMRG value, we plot in the inset of Fig. 1b the minimum number of parameters that is required in order to reach an error value $\Delta=\frac{E_{\text{VQE}}-E_{\text{DMRG}}}{N}$ of $0.01$ with the NP ansatz and for 2D lattices. We see that the necessary number of parameters to reach $\Delta=0.01$ scales as $(N_{x}\times N_{y})^{n}$, with $n=2.13$. A similar power-law behavior with an exponent $n$ in the range of $1.6-2$ is found for 1D systems depending on the value of $\Delta$.
@@
We now turn to results obtained with the EP ansatz. Figs. 1c and 1d visualize the error per lattice site of VQE with respect to DMRG for $U/t=2$ and for one- and two-dimensional lattices, respectively. As for the case of the NP ansatz, we optimize PQCs of up to thirteen layers. For one-dimensional lattices, the performance of VQE based on the EP ansatz is very similar to that of VQE based on the NP ansatz. However, for two-dimensional lattices, it is immediately evident that the VQE energy converges to the DMRG one at a much slower rate than when using the NP ansatz. Moreover, for larger lattices we observe that the VQE energy plateaus at substantial error values; when using $13$ layers of this ansatz for a $3\times 3$ lattice for example, the error per lattice site of the VQE energy compared to the DMRG value is $0.019$-$0.029$, compared to $0.0003$-$0.001$ for the NP ansatz. We attribute this behavior to the fact that the EP ansatz does not directly couple vertically adjacent qubits in two-dimensional lattices, due to the lack of fermionic swaps, hence missing a key part of the physics of the Hubbard model in which entanglement is generated efficently and directly between neighboring lattice sites.
@@
The UCC ansatz is a popular choice for VQE calculations, particularly for chemistry applications. Given this popularity, we benchmark in Fig. 2 the performance of the NP, EP and UCC ansätze for the solution of a $2\times 3$ Hubbard model with $U/t=2$. Unlike the layered NP and EP ansätze, the UCC ansatz contains a set number of one- and two-electron excitation operators, and a number of parameters that is determined only by the lattice size. We see in Fig. 2 that for our case study, the UCC ansatz requires significantly more parameters (quadratic scaling, see Section II.2) compared to NP or EP PQCs of up to thirteen layers (linear scaling), thus leading to a higher computational cost. Indeed we were not able to study lattices larger than $2\times 3$ using the UCC ansatz due to the steep increase in the number of optimization parameters - for a $3\times 3$ lattice, the ansatz has $3213$ parameters, more than three times that of the NP ansatz at thirteen layers. Moreover, we see in Fig. 2 that despite the larger number of optimization parameters associated with the UCC ansatz, there is no significant improvement of the results compared to the EP ansatz, while the NP ansatz significantly outperforms it. In the future, it will be worthwhile to explore ways to optimize the UCC ansatz for lattice model simulations such as by adaptive approaches or by truncating excitation operators through a ranking scheme of their relative importance, which has been implemented for molecular systems through an MP2 pre-optimization scheme [56]. Another striking observation when comparing the layered ansätze is that, for small parameter numbers $\lesssim 150$, the EP ansatz actually slightly outperforms NP. However the rate of convergence with the number of parameters slows down markedly for the EP ansatz as this parameter number threshold is surpassed, and the NP ansatz systematically yields smaller errors. We find that it is a general feature across different lattices that for smaller number of parameters the NP and EP ansätze perform similarly, however the energy error obtained from optimizing the EP PQC plateaus for larger numbers of parameters.
@@
III.2 Impact of interaction strength $U/t$
Figure 4: Comparison of the energy difference per lattice site, of the VQE ground state from DMRG for 2D lattices, when using the NP ansatz and for $U/t=2$, in the presence of disorder $d$ (panel a) and off-site Coulomb interactions $V$ (panel b). Solid lines represent the case of no disorder/off-site Coulomb interaction $d=0,V=0$, dotted lines represent $d=0.2$ and $V=0.2$ respectively, while dashed lines represent $d=0.8$ and $V=0.8$ respectively.
Up to this point, we have focused on the case with interaction strength $U/t=2$. However, for several applications of interest in condensed matter physics, stronger electronic correlations become important, and previous results on $1\times 8$ Hubbard chains indicate that this regime might be more challenging to capture with VQAs [40]. To address the quality of VQE simulations in this strong coupling regime, we set $U/t=8$. Motivated by the results of Fig. 1, we adopt the NP ansatz in the following. Figs. 3a and 3b demonstrate the convergence of the VQE results towards the DMRG ones as a function of the number of parameters in the PQC for various lattice sizes in 1D and 2D, respectively. It becomes clear from these results that stronger electronic correlations pose more of a challenge to accurate VQE simulations of the Hubbard model, even when using the NP ansatz. To emphasize this, we plot in Fig. 3c the VQE energy error per site for 2D lattices, for $U/t=2$ and $U/t=8$. It is also worth noting that similar to $U/t=2$ and Fig. 1b, we find for $U/t=8$ that there is a power-law dependence of the minimum number of parameters needed to achieve an accuracy of $\Delta=0.01$, on the lattice size, with an exponent of $n=2.18$ for 2D systems and $n=1.61$ for 1D systems.
@@ III.2 Impact of interaction strength $U/t$
III.3 Impact of disorder and off-site Coulomb interactions
An important factor towards increasingly realistic materials simulations on quantum hardware is spatial inhomogeneity not present in the simple Hubbard model of Eq. (1). Indeed, the material-specific Hubbard models of the form (I), derived through methodologies such as ab initio downfolding [23, 24, 25] and embedding  [26, 27], allow for spatial variations in the hopping and interaction parameters, e.g., due to static disorder present in the material.
Figure 5: Comparison of the energy difference per lattice site, of the VQE ground state from DMRG, as obtained from an energy-based optimization (solid lines) and an overlap-based optimization (dashed lines), when using the NP ansatz. The different panels visualize the results for 1D and 2D lattices, and for the cases $U/t=2$ and $U/t=8$.
Figure 6: Comparison of the fidelity of the VQE ground state with respect to DMRG, as obtained from an energy-based optimization (solid lines) and an overlap-based optimization (dashed lines), when using the NP ansatz, and for 2D lattices with $U/t=2$ (panel a) and $U/t=8$ (panel b).
@@ III.3 Impact of disorder and off-site Coulomb interactions
where we now have disorder in the nearest-neighbor hopping ($\mathbf{R}\neq\mathbf{R}^{\prime}$) and chemical potential ($\mathbf{R}=\mathbf{R}^{\prime}$) terms, $t_{\mathbf{R}\mathbf{R}^{\prime}}=t\cdot\delta_{\langle\mathbf{R}\mathbf{R}^{\prime}\rangle}+\delta t_{\mathbf{R}\mathbf{R}^{\prime}}$. The function $\delta_{\langle\mathbf{R}\mathbf{R}^{\prime}\rangle}$ is equal to one for nearest neighbors, and zero otherwise, since the chemical potential in our system is set to zero. The disorder enters via $\delta t_{\mathbf{R}\mathbf{R}^{\prime}}=d\cdot\mathcal{N}(0,1)$, with $d$ a user-defined parameter for the magnitude of the disorder and $\mathcal{N}(0,1)$ a function drawing a random number from a Gaussian distribution of zero mean and unit variance. In this way we ensure that the values of $t_{\mathbf{R}\mathbf{R}^{\prime}}$ of nearest neighbor hopping are centered around $t$, and the chemical potential is centered around zero, both with a standard deviation of $d$. In Fig. 4a we plot the error per lattice site of the VQE ground state energies with respect to the DMRG ones, for different lattices and as a function of the number of optimization parameters in the ansatz. We use disorder values of $d=0.2$ and $d=0.8$. We observe a small increase in the error of the VQE energies as compared to the homogeneous case ($d=0$), however this increase is not monotonic with the value of $d$. Despite this small increase in the VQE energy error, these results suggest that even for disordered systems with nonuniform parameters, VQE using the NP ansatz can recover the ground state of the Hubbard model with reasonable accuracy, within different regimes of disorder.
So far we have focused on Hubbard models where Coulomb interactions are present only on-site. However, in realistic material systems there are also longer-range interactions present, which motivates us to investigate the impact of including a finite nearest-neighbor repulsive Coulomb interaction $V$ in our model, i.e., to consider models of the form
@@
In Fig. 4b we visualize the energy error per lattice site of the VQE solutions of the Hubbard model on 2D lattices, with respect to the DMRG energy, as a function of the number of parameters used in the optimization, and for different values of $V$, specifically $V=0$ (solid lines), $V=0.2$ (dotted lines) and $V=0.8$ (dashed lines). Similar to the case of Fig. 4a of having finite disorder in the lattice, we see that while indeed having off-site interactions generally leads to slightly larger errors in the VQE energy, the increase of the error is non-monotonic with $V$. The fact that VQE with the NP ansatz seems capable of accurately describing spatially inhomogeneous systems with Coulomb interactions beyond on-site is encouraging for the simulation of increasingly complex and inhomogeneous systems on quantum hardware. We relate this capability to the fact that, unlike in the original HVA ansatz [33], the NP ansatz contains independent variational parameters on every link that appears in the Hamiltonian.
III.4 Impact of using an overlap-based optimization
Here we compare the performance of conventional VQE, where we drive the optimization by minimizing the energy expectation value of the wavefunction, to overlap-based optimization, where we minimize the logarithm of the infidelity with respect to the DMRG ground state (eq. (5)). In Fig. 5 we visualize the energy error per lattice site of VQE based on overlap optimization (dashed lines), alongside the values previously obtained from energy optimization (solid lines), for 1D and 2D lattices, and for the cases of $U/t=2$ and $U/t=8$. We find that in the one-dimensional case the overlap-based optimization leads to a small improvement of the VQE energies, for both $U/t=2$ and $U/t=8$. This suggests that the residual energy error for larger lattices is a result of the complex cost function landscape rather than the expressivity of the NP ansatz. However, for two-dimensional lattices we find no substantial differences between the two methods, and the energy-based optimization can even slightly outperform the overlap-based one in some cases.
It is also worth investigating how the energy- and overlap-based optimization perform in terms of producing high-fidelity VQE ground states. Figs. 6a and 6b visualize the fidelity $\mathcal{F}=|\bra{\Psi_{\rm VQE}}\ket{\Psi_{\rm DMRG}}|^{2}$ of the VQE ground state, when using the NP ansatz for 2D lattices, and for $U/t=2$ and $U/t=8$. We see that with the addition of more parameters to the optimization, both methods converge towards the limit of $\mathcal{F}=1$, except for the largest system size $3\times 3$. However, the convergence of the overlap-based optimization is generally faster, especially for larger lattices, a feature which is particularly prominent in the case of $U/t=8$, where strong electronic correlations are present. The fact that the overlap-based optimization performs better in terms of yielding higher-fidelity ground states is not surprising, given that this method is tailored specifically to minimize the loss function of eq. (5), i.e., to minimize the infidelity of the VQE ground state. However, it is worth emphasizing that the higher fidelity solutions produced via overlap-based optimization do not always possess lower energies compared to the lower-fidelity wavefunctions obtained via energy-based minimization. An example of this is seen for $3\times 3$ lattices, and for both $U/t=2$ and $U/t=8$. While overlap-based optimization produces ground states of higher fidelities as seen in Fig. 6, the energy error of these states is in fact larger than that from energy-based optimization, as seen in Fig. 5.
@@ III.4 Impact of using an overlap-based optimization
III.5 Scaling up and extracting correlation functions
Our results so far have demonstrated that convergence of the VQE energy to the DMRG value with respect to the number of the optimization parameters becomes slower for larger lattice sizes. It is therefore interesting to explore this convergence in more detail for a larger lattice of size $4\times 4$ for $U/t=2$, to examine whether this trend holds even for larger systems. Since each lattice site can be occupied by a spin-up and/or a spin-down electron, this corresponds to $4\times 4\times 2=32$ qubits. Due to the large computational cost associated with simulating a lattice of this size, we restrict the bond dimension of the MPS representation within our VTNE approach, and also within DMRG, to $\chi=512$. Here we also set $d=V=0$. As in previous cases, we initialize ten VQE optimizations from different, random starting points. Two of these failed to converge, and for the remaining eight we visualize their converged energy errors per lattice site with respect to DMRG in Fig. 7. The lowest energy obtained through VQE optimization, with 12 NP ansatz layers corresponding to 1,568 parameters, has a relative error of $2.2\%$ in its energy, compared to the DMRG value. This demonstrates that even for these larger lattices, reasonable accuracy for the ground state of the Hubbard model can be achieved using VQE, as long as enough parameters are included in the optimization.

### v2

@@ 2.2.3 Unitary coupled cluster (UCC) ansatz
3 Results
Figure 3: Energy difference per lattice site, between VQE and DMRG ground state energies when using the NP and EP ansätze, for 1D and 2D lattices, on a logarithmic scale. The shaded region indicates the difference between the minimum and maximum ground state energies obtained using VQE with a total of ten different random starting points. The inset of panel a shows a log-log plot of the minimum number of parameters needed to achieve an error per lattice site $\Delta=0.01$, when using the NP ansatz for 1D lattices.
@@ 3 Results
We now systematically study the impact of varying the number of variational parameters and the ansatz, the strength of the electronic interactions, and the strength of disorder and off-site Coulomb interactions, on the quality of the VQE solutions of the half-filled Hubbard model with respect to DMRG. Moreover, we benchmark the performance of using an overlap-based optimization, compared to the more conventional energy-based optimization. The optimized VQE energies for all cases presented below are given in the Supplemetary Material.
3.1 Impact of the ansatz and the number of variational parameters
@@
We start our analysis by describing the VQE result for the half-filled Hubbard model of different system sizes in 1D and 2D square lattice geometries. We employ open boundary conditions throughout. We will first explore how the NP ansatz, which was designed specifically for solving the Hubbard model on quantum hardware, performs as a function of lattice size and number of ansatz parameters. Our benchmarks make use of PQCs of up to thirteen layers.
In Fig. 3a we plot for several one-dimensional lattices the difference between our VQE energy and the DMRG benchmark per lattice site, for $U/t=2$. This interaction strength lies in the weakly- to moderately-correlated regime [67, 34]. We note that in the thermodynamic limit both 1D and 2D Hubbard models for $U>0$ are in a Mott insulating phase at half filling with dominant antiferromagnetic Néel-type correlations [36, 68]. While the charge gap is nonzero, the spin gap vanishes. As our ground state energy from VQE we plot our best estimate, obtained as the lowest value from ten independent optimizations with random starting points. The shaded regions in Fig. 3a indicate the full range of energies obtained over all optimizations. In Fig. 3a we only visualize the results for even numbers of sites, however, in the Supplemental Material we also provide the results for odd lattices, which are consistent with the trends observed here.
@@
While it is clear that the VQE optimization converges to the ground state energy even for the larger lattices as more parameters are added to the PQC, the convergence becomes increasingly slow for larger lattice sizes, and even plateaus to errors of order $10^{-2}$ in those cases. For the $1\times 12$ lattice with $13$ layers of the NP ansatz we have also performed an extended optimization, where we performed a hundred independent optimizations (rather than ten) starting from different random parameters. Nevertheless, even when we explore a larger part of phase space, we do not find any appreciable reduction in the ground state energy produced by VQE. Fig. 3b visualizes the performance of VQE on a logarithmic scale for two-dimensional lattices at $U/t=2$. We again see that the convergence of VQE using the NP ansatz slows down with increasing system size, consistent with our benchmarks of the one-dimensional case. This plateauing is the result of the complex optimization landscape of larger systems with a significant number of variational parameters, which makes VQE prone to becoming stuck in local minima. This suggests further work could be devoted to using a combination of optimizers including genetic algorithms. While the energy gradients during the course of our optimizations generally always remain finite and we do not encounter barren plateaus, these could appear and pose a challenge on quantum hardware, where the numerical accuracy of gradients is limited by shot noise.
@@
To further understand the convergence behavior of the VQE energy towards the DMRG value, we plot in the inset of Fig. 3a the minimum number of parameters that is required in order to reach an error value $\Delta=\frac{E_{\text{VQE}}-E_{\text{DMRG}}}{N}$ of $0.01$ with the NP ansatz and for 1D lattices. We see that the necessary number of parameters to reach $\Delta=0.01$ scales as $(N_{x}\times N_{y})^{n}$, with $n=1.61$. A similar power-law behavior with an exponent $n$ of $2.13$ is found for 2D systems, when we aim to reach the same error value. It should be emphasized that this power-law behavior is established empirically through our analysis of the numerical results we obtain in this work, nevertheless, it will be useful for anticipating the computational cost of VQE simulations for larger systems.
@@
We now turn to results obtained with the EP ansatz. Figs. 3c and 3d visualize the error per lattice site of VQE with respect to DMRG for $U/t=2$ and for one- and two-dimensional lattices, respectively. Once again, for visualization purposes we only show the results for even one-dimensional lattices, and include all data for odd lattices in the Supplemental Material. As for the case of the NP ansatz, we optimize PQCs of up to thirteen layers. In all cases, the performance of VQE based on the EP ansatz is worse than that of VQE based on the NP ansatz. As already discussed, this is unsurprising given the generic character of the EP ansatz, compared to the NP ansatz where different ansatz components represent specific physical terms of the Hamiltonian. The difference between the two becomes particularly pronounced for two-dimensional lattices, where it is immediately evident that the VQE energy converges to the DMRG one at a much slower rate than when using the NP ansatz. Moreover, for larger lattices we observe that the VQE energy plateaus at substantial error values; when using $13$ layers of this ansatz for a $3\times 3$ lattice for example, the error per lattice site of the VQE energy compared to the DMRG value is $0.019$-$0.029$, compared to $0.0003$-$0.001$ for the NP ansatz. We attribute this behavior to the fact that the EP ansatz does not directly couple vertically adjacent qubits in two-dimensional lattices, due to the lack of fermionic swaps, hence missing additional key parts of the physics of the Hubbard model. We conclude that the performance of EP is limited by its expressivity. Another interesting observation when comparing the layered ansätze is that, for small parameter numbers $\lesssim 150$, the EP ansatz performs similar to NP. However, the rate of convergence with the number of parameters slows down markedly for the EP ansatz as this parameter number threshold is surpassed, and the NP ansatz systematically yields smaller errors. We find that it is a general feature across different lattices that for smaller number of parameters the NP and EP ansätze perform similarly, however the energy error obtained from optimizing the EP PQC plateaus for larger numbers of parameters.
@@
The UCC ansatz is a popular choice for VQE calculations, particularly for chemistry applications. Given this popularity, we benchmark in Fig. 4 the performance of the NP, EP and UCC ansätze for the solution of a $2\times 3$ Hubbard model with $U/t=2$ (here we only show the lowest VQE energy obtained for each number of parameters and for each ansatz). Unlike the layered NP and EP ansätze, the UCC ansatz contains a set number of one- and two-electron excitation operators, and a number of parameters that is determined only by the lattice size. We see in Fig. 4 that for our case study, the UCC ansatz requires significantly more parameters (quadratic scaling, see Section 2.2) compared to NP or EP PQCs of up to thirteen layers (linear scaling), thus leading to a higher computational cost. Indeed we were not able to study lattices larger than $2\times 3$ using the UCC ansatz due to the steep increase in the number of optimization parameters - for a $3\times 3$ lattice, the ansatz has $3213$ parameters, more than three times that of the NP ansatz at thirteen layers. Moreover, we see in Fig. 4 that despite the larger number of optimization parameters associated with the UCC ansatz, there is no significant improvement of the results compared to the EP ansatz, while the NP ansatz significantly outperforms it. In the future, it will be worthwhile to explore ways to optimize the UCC ansatz for lattice model simulations such as by adaptive approaches or by truncating excitation operators through a ranking scheme of their relative importance, which has been implemented for molecular systems through an MP2 pre-optimization scheme [69].
Our results so far have demonstrated that convergence of the VQE energy to the DMRG value with respect to the number of the optimization parameters becomes slower for larger lattice sizes. It is therefore interesting to explore the convergence of the ground state energy in more detail for a larger lattice of size $4\times 4$ for $U/t=2$, using the NP ansatz. Since each lattice site can be occupied by a spin-up and/or a spin-down electron, this corresponds to $4\times 4\times 2=32$ qubits. Due to the large computational cost associated with simulating a lattice of this size, we restrict the bond dimension of the MPS representation within our VTNE approach, and also within DMRG, to $\chi=512$. In the Supplemental Material we demonstrate the convergence of the DMRG energy per lattice site with the bond dimension, which allows us to estimate that using $\chi=512$ leads to an error of approximately $5\times 10^{-3}$ for this quantity. As in previous cases, we initialize ten VQE optimizations from different, random starting points. Two of these failed to converge, and for the remaining eight we visualize their converged energy errors per lattice site with respect to DMRG in Fig. 5. We see that the energy errors have a wide distribution, indicative of the fact that some optimizations become stuck in local minima. Nevertheless, in some cases we can approach the ground state energy with reasonable accuracy. To reach this level of accuracy, we allowed each optimization to run for $48$ hours on a single CPU of NERSC Perlmutter, and each utilized an average of $7.1$ GBs of memory. The lowest energy obtained through VQE optimization, with 12 NP ansatz layers corresponding to 1,568 parameters, has a relative error of $2.2\%$ in its energy, compared to the DMRG value. This demonstrates that even for these larger lattices, VQE using the NP ansatz can, at least in principle, yield ground states of the Hubbard model that are not far from the ground truth, as long as enough parameters are included in the optimization. We conclude that the NP ansatz is sufficiently expressive, but the optimization is challenging due to local minima.
@@
3.2 Impact of interaction strength $U/t$
Figure 7: Comparison of the energy difference per site of a $3\times 3$ lattice, of the VQE ground state from DMRG, when using the NP ansatz and for $U/t=2$, in the presence of disorder $d$ (panel a) and nearest-neighbor (screened) Coulomb interactions $V$ (panel b). The shading represents the range of energies obtained in the different cases.
Up to this point, we have focused on the weakly-correlated case with interaction strength $U/t=2$. However, for several applications of interest in condensed matter physics, stronger electronic correlations become important, and previous results on $1\times 8$ Hubbard chains indicate that this regime might be more challenging to capture with VQAs [46]. To address the quality of VQE simulations in this strong coupling regime, we set $U/t=8$. Motivated by the results of Fig. 3, we adopt the NP ansatz in the following. Figs. 6a and 6b demonstrate the convergence of the VQE results towards the DMRG ones as a function of the number of parameters in the PQC for various lattice sizes in 1D and 2D, respectively. It becomes clear from these results and by comparing them to those of Fig. 3 that stronger electronic correlations pose more of a challenge to accurate VQE simulations of the Hubbard model, even when using the NP ansatz. To emphasize this, we visualize in Fig. 6c the VQE energy error per site for $2\times 3$ and $3\times 3$ lattices, for $U/t=2$ (solid lines) and $U/t=8$ (dashed lines) on the same plot. It is also worth noting that similar to $U/t=2$ and Fig. 3b, we find for $U/t=8$ that there is a power-law dependence of the minimum number of parameters needed to achieve an accuracy of $\Delta=0.01$, on the lattice size, with an exponent of $n=2.18$ for 2D systems and $n=1.61$ for 1D systems.
@@ 3.2 Impact of interaction strength $U/t$
3.3 Impact of disorder and off-site Coulomb interactions
An important factor towards increasingly realistic materials simulations on quantum hardware is spatial inhomogeneity not present in the simple Hubbard model of Eq. (1). Indeed, the material-specific Hubbard models of the form of Eq. (1), derived through methodologies such as ab initio downfolding [25, 26, 27, 50] and embedding  [28, 29], allow for spatial variations in the hopping and interaction parameters, e.g., due to the presence of different chemical elements or of static disorder within the material.
@@ 3.3 Impact of disorder and off-site Coulomb interactions
where we now have disorder in the nearest-neighbor hopping ($\mathbf{R}\neq\mathbf{R}^{\prime}$) and on-site potential ($\mathbf{R}=\mathbf{R}^{\prime}$) terms, $t_{\mathbf{R}\mathbf{R}^{\prime}}=t\cdot\delta_{\langle\mathbf{R}\mathbf{R}^{\prime}\rangle}+\delta t_{\mathbf{R}\mathbf{R}^{\prime}}$. The function $\delta_{\langle\mathbf{R}\mathbf{R}^{\prime}\rangle}$ is equal to one for nearest neighbors, and zero otherwise, since the on-site potential in our system is set to zero. The disorder enters via $\delta t_{\mathbf{R}\mathbf{R}^{\prime}}=d\cdot\mathcal{N}(0,1)$, with $d$ a user-defined parameter for the magnitude of the disorder and $\mathcal{N}(0,1)$ a function drawing a random number from a Gaussian distribution of zero mean and unit variance. The values of $t_{\mathbf{R}\mathbf{R}^{\prime}}$ for nearest neighbor hopping are centered around $t$, and the on-site potential is centered around zero, both with a standard deviation of $d$. In Fig. 7a we plot the error per lattice site of the VQE ground state energies with respect to the DMRG ones, for a $3\times 3$ lattice, and as a function of the number of optimization parameters in the ansatz. We use disorder values of $d=0.2$ (red) and $d=0.8$ (blue). The results suggest that even for disordered systems with nonuniform parameters, VQE using the NP ansatz can recover the ground state of the Hubbard model with similar accuracy to the disorder-free case $d=0$ (gray). We make similar findings irrespective of the lattice size, and we do not include results beyond the $3\times 3$ case in Fig. 7 for visibility purposes. All data for disordered simulations on different lattices are given in the Supplementary Material. Interestingly, we observe in Fig. 7 that the introduction of disorder, particularly for $d=0.8$, actually reduces the error of the VQE energy with respect to the DMRG value. One reason behind this could be that the symmetry-breaking effect of disorder lifts degeneracies, effectively simplifying the VQE optimization landscape. Moreover, disorder induces localization, which can reduce entanglement [71], hence making it easier to capture the ground state with shallower circuits.
So far we have focused on Hubbard models where Coulomb interactions are only present for electrons on the same site. However, in realistic material systems there are also longer-range interactions present, which motivates us to investigate the impact of including a finite nearest-neighbor repulsive (screened) Coulomb interaction $V$ in our model, i.e., to consider extended Hubbard models of the form
@@
In Fig. 7b we visualize the energy error per lattice site of the VQE solutions of the extended Hubbard model on a $3\times 3$ lattice, with respect to the DMRG energy, as a function of the number of parameters used in the optimization, and for different values of $V$, specifically $V=0$ (gray), $V=0.2$ (red), and $V=0.8$ (blue). Similar to the case of Fig. 7a of having finite disorder in the lattice, we see that having nearest-neighbor interactions generally leads to comparable errors in the VQE energy to when $V=0$. We make similar findings for all lattice sizes studied here, with all numerical results given in the Supplementary Material. The fact that VQE with the NP ansatz seems capable of describing generalized, extended Hubbard models that include spatial inhomogeneity and (screened) Coulomb interactions beyond on-site, with a similar degree of accuracy as for the standard Hubbard model of Eq. (1), is encouraging for the simulation of increasingly complex, realistic and inhomogeneous systems on quantum hardware. We relate this capability to the fact that, unlike in the original HVA ansatz [38], the NP ansatz contains independent variational parameters on every bond that appears in the Hamiltonian.
@@
3.4 Extracting correlation functions and overlap-based optimization